Compute the 2018 tax liability and the marginal and average tax rates for the following taxpayers. Click here to access the 2018 tax rate schedule. If required, round the tax liability the nearest dollar. When required, round the average rates to four decimal places before converting to a percentage (i.e. .67073 would be rounded to .6707 and entered as 67.07%). a. Chandler, who files as a single taxpayer, has taxable income of $100,000.

Tax liability: $

Marginal rate: %

Average rate: %

b. Lazare, who files as a head of household, has taxable income of $72,200.

Tax liability:

$ Marginal rate: %

Average rate: %
Answer #1

$100,000; single:

The taxable income fall under 24% tax bracket; therefore, the marginal tax rate is 24%.

The corresponding figures are shown in the table under such bracket.

The tax liability = $14,089.50 + (100,000 – 82,500) × 24%

                        = $14,089.50 + $17,500 × 24%

                        = $14,089.50 + $4,200

                        = $18,289.50 (Answer)

Average tax rate = (Tax liability / Taxable income) × 100

                           = ($18,289.50 / $100,000) × 100

                           = (0.182895) × 100

                           = 0.1829 × 100

                           = 18.29%

Answers:

Tax liability = $18,289.50

Marginal tax rate = 24%

Average tax rate = 18.29%

$72,200; head of household:

The taxable income fall under 22% tax bracket; therefore, the marginal tax rate is 22%.

The corresponding figures are shown in the table under such bracket.

The tax liability = $5,944 + (72,200 – 51,800) × 22%

                        = $5,944 + $20,400 × 22%

                        = $5,944 + $4,488

                        = $10,432 (Answer)

Average tax rate = (Tax liability / Taxable income) × 100

                           = ($10,432 / $72,200) × 100

                           = (0.144487…..) × 100

                        = 0.1445 × 100

                           = 14.45%

Answers:

Tax liability = $10,432

Marginal tax rate = 22%

Average tax rate = 14.45%